# DSR-relativistic spacetime picture and the phenomenology of Planck-scale-modified time dilation

### v1

@@ I Preliminaries on time dilation and dispersion relation
As we already observed, in the LIV scenario time dilation is automatically unmodified; therefore, our last task is to investigate time dilation in a DSR scenario with dispersion relation (1), and this is a rather technical challenge for which the previous DSR literature provides no insight. A large number of results has been obtained on DSR-relativistic frameworks, but the time dilation was never studied. This is essentially because, as we will see, both the finite boost transformations on coordinates and the full invariant spacetime metric were missing in the literature of DSR. The rest of this manuscript is devoted to fill this gap.
We will go beyond the standard first-order approximation typical of DSR studies, working at all orders of the deformation parameter $\ell$. This will allow us to establish several DSR properties that were previously unknown or only partially understood.
In Section II, we introduce the specific DSR scenario on which we base our analysis. In Section III, we derive finite boost transformations of all phase-space variables. In Section IV, we establish properties of the worldlines that are needed for the derivation of time dilation. In Section V, we determine the DSR-relativistic spacetime metric. In Section VI, we derive the time dilation factor. The final Section is devoted to a summary of our results and the analysis of its phenomenological implications.
@@ III Finite boost transformation of phase-space variables
After substituting the solution (11) in the equations (10) for spacetime coordinates and setting initial conditions $\overline{x}^{0}$$=$$x^{0}(0)$ and $\overline{x}^{1}$$=$$x^{1}(0)$, we can solve the other two equations in (10), describing the transformation of spacetime coordinates, and find
@@
We start by studying their domain of existence. In special relativity, the rapidity $\eta$ can take any real value. In this DSR model, instead, such property holds only for $\ell>0$, while for negative $\ell$ the solutions are valid within the finite range $\eta_{-}$$<$ $\eta<\eta_{+}$, where
@@ IV Worldline covariance
where $x^{\prime}(0)\equiv x^{\prime}_{\epsilon^{\prime},p^{\prime}}(t^{\prime}$$=$0)0)0 ). Concerning velocity, a simple computation shows that $V^{\prime}=V(\epsilon^{\prime},p^{\prime})$ where $V^{\prime}\equiv\frac{dx^{\prime}}{dt^{\prime}}$ is read from Eq. (25) and the functional form of $V(\epsilon^{\prime},p^{\prime})$ is the one of Eq. (22), with $\epsilon^{\prime}$ and $p^{\prime}$ given by Eq. (24). Notice that
@@
The results obtained in this Section so far allow us to find the relation between a particle’s velocity $V$ and rapidity $\eta$. In special relativity, they are connected by the familiar relation $V=\tanh\eta$. In DSR, this standard relation is deformed. To see this, let us consider a particle at rest, so that $\epsilon^{\prime}=M$, $p^{\prime}=0$, and $V^{\prime}=0$. After a boost with rapidity $-\eta$, the particle’s velocity is given by (27), which in this case reads $V=\Lambda_{10}(M,0;-\eta)/\Lambda_{00}(M,0;-\eta)$. Writing the matrix elements of $\Lambda(M,0;-\eta)$ explicitly we find the deformed relation
@@ V Covariant spacetime metric
It is well established that the mass-shell relation of a DSR theory defines a metric on momentum space, such that the mass-shell relation is given by the geodesic distance of the point $(\epsilon,p)$ from the origin [32, 33, 34]. [^5] Other properties of DRS models, such as the laws of addition of energy and momentum, define torsion and nonmetricity [33, 37] Because of the relativistic properties of DSR models, these possibly curved momentum space metrics are maximally symmetric, that is they allow for a full set of relativistic generators. For instance, the momentum space metric of the DSR model we are considering here is known to be that of a de Sitter manifold [35, 36, 33, 34, 37, 38, 39]: the mass-shell relation (7) can be obtained from the metric $\tilde{g}(\epsilon)=\text{diag}(1,-e^{2\ell\epsilon})$, which is indeed a de Sitter metric on momentum space. [^6] An example of anti-de Sitter momentum spaces can be found in [40].
@@ VI Time Dilation
In special relativity, the proper time interval $\Delta\tau$ of a particle and the corresponding time interval $\Delta t$ measured by a boosted observer are related by $\Delta t=\gamma_{sr}\Delta\tau$, where the time dilation factor is $\gamma_{sr}$$=\cosh{\eta}=\epsilon/M$, written in terms of rapidity $\eta$ or of energy $\epsilon$ of the particle, respectively. This special-relativistic result can be equivalently obtained either from the Lorentz transformation of the time interval or from the Minkowski metric. In this Section, we use the relations derived so far to compute the time dilation factor in the DSR model, showing that the two methods give equivalent (but deformed) results also in the DSR scenario.
@@
[…]
@@
The time dilation factor $\gamma_{dsr}$ reduces to the standard factor $\gamma_{sr}$$=\cosh{\eta}=\epsilon/M$ in the undeformed limit $\ell\to 0$ and is fully analytical in the deformation parameter $\ell$, in contrast to previous proposals described in Section I. Moreover, $\gamma_{dsr}$ is well-defined, positive and monotonous for all possible values of $M$ and $\epsilon$ (or $\eta$), that is, it has no pathologies and always preserves causality. Looking at the infinite-boost limit, when $\ell>0$ the time dilation factor $\gamma_{dsr}$ saturates to the finite value $\coth{(\ell M)}$ for $\eta\to\pm\infty$, while for negative deformation parameter $\ell<0$ the infinite boost limit corresponds to $\eta\to\eta_{\pm}$ and the factor $\gamma_{dsr}$ is divergent. [^8] The infinite-boost limit can be equivalently taken in terms of energy: for both signs of $\ell$ it reads $\epsilon\to\infty$ and one finds the same saturation property of $\gamma_{dsr}$, as can be directly checked from (41). This can be compared to special relativity, where the time dilation factor $\gamma_{sr}$ diverges for $\eta\to\pm\infty$.
@@ VII Summary and Outlook
In this paper, we investigated the time dilation effect within theories characterized by a Planck-scale modified dispersion relation of the form (1). Having observed that previous proposals, described by Eq. (2) and derived independently in [13] and [15], are consistent neither with standard Lorentz-invariance violation (LIV) scenarios nor with DSR scenarios, we derived the time dilation effect within a fully DSR-relativistic model compatible with the dispersion relation (1). To accomplish this task, we needed to derive some properties of a DSR spacetime picture, which were not available in the literature.
We derived the finite boost transformations on all phase-space variables and to all orders in the deformation parameter. We showed that free particle worldlines are fully covariant under these boost transformations, generalizing previous results obtained at the first order in the deformation parameter, and we identified the covariant spacetime metric. These results provide the ingredients needed to compute the time dilation factor, encoded in Eq.(40). Having obtained a formula valid to all orders in the deformation parameter, we could verify that the associated dilation factor $\gamma_{dsr}$ is well-defined and monotonous for all boosts. This is in contrast with other descriptions of time dilation beyond special relativity, encoded in Eq. (2), as discussed in Section I. This qualitative difference translates into different phenomenological implications.
@@
Not surprisingly, the correction with respect to the standard factor is governed by $\ell\epsilon$, a quantity that is small as long as the energy is subplanckian. In other words, our fully relativistic DSR derivation provides an analytical deformation of the standard time dilation formula. Because of this, phenomenological effects that were estimated to be significant for the scenarios leading to Eq. (2) are much smaller in this scenario.

### v2

@@ I Preliminaries on time dilation and dispersion relation
As we already observed, in the LIV scenario time dilation is automatically unmodified; therefore, our last task is to investigate time dilation in a DSR scenario with dispersion relation (1). And for this task the literature provides only partial and/or indirect supporting material. Several alternative DSR scenarios have been considered, with or without upholding (1) (see, e.g., Refs [20, 19]), and the focus of the analyses is typically oriented toward momentum-space features rather than spacetime observables. DSR-relativistic spacetime observables were studied in Refs [21, 22], relying however on spacetime coordinates governed by non-standard Poisson brackets, whose relation with standard coordinates suitable for describing the relativistic properties of particles in a laboratory setup remains unclear. We here investigate time dilation for the most studied DSR scenario (reviewed briefly in the next section), which indeed upholds (1), and, in order to allow direct comparison with the results of [13, 14, 15], we work with spacetime coordinates that satisfy standard (i.e. trivial) Poisson brackets. After devoting the next Section to the introduction of the specific DSR scenario on which we base our analysis, in Section III, we derive finite boost transformations of all phase-space variables. In Section IV, we establish properties of the worldlines that are needed for the derivation of time dilation. In Section V, we determine the DSR-relativistic spacetime metric. In Section VI, we derive the time dilation factor. In the closing Section VII we summarize our main results and discuss their phenomenological implications.
@@ III Finite boost transformation of phase-space variables
After substituting the solution (11) in the equations (10) for spacetime coordinates and setting initial conditions $\overline{x}^{0}$$x^{0}(0)$ and $\overline{x}^{1}$$x^{1}(0)$, we can solve the other two equations in (10), describing the transformation of spacetime coordinates, and find
@@
We start by studying their domain of existence. In special relativity, the rapidity $\eta$ can take any real value. In this DSR model, instead, such property holds only for $\ell>0$, while for negative $\ell$ the solutions are valid within the finite range $\eta_{-}$<< $\eta<\eta_{+}$, where
@@ IV Worldline covariance
where $x^{\prime}(0)\equiv x^{\prime}_{\epsilon^{\prime},p^{\prime}}(t^{\prime}$==$0)$. Concerning velocity, a simple computation shows that $V^{\prime}=V(\epsilon^{\prime},p^{\prime})$ where $V^{\prime}\equiv\frac{dx^{\prime}}{dt^{\prime}}$ is read from Eq. (25) and the functional form of $V(\epsilon^{\prime},p^{\prime})$ is the one of Eq. (22), with $\epsilon^{\prime}$ and $p^{\prime}$ given by Eq. (24). Notice that
@@
The results obtained in this Section so far allow us to find the relation between a particle’s velocity $V$ and rapidity $\eta$. In special relativity, they are connected by the familiar relation $V=\tanh\eta$. In DSR, this standard relation is deformed. To see this, let us consider a particle at rest, so that $\epsilon^{\prime}=M$, $p^{\prime}=0$, and $V^{\prime}=0$. After a boost with rapidity $\eta$, the particle’s velocity is given by (27), which in this case reads $V=\Lambda_{10}(M,0;-\eta)/\Lambda_{00}(M,0;-\eta)$. Writing the matrix elements of $\Lambda(M,0;-\eta)$ explicitly we find the deformed relation
@@ V Covariant spacetime metric
It is well established that the mass-shell relation of a DSR theory defines a metric on momentum space, such that the mass-shell relation is given by the geodesic distance of the point $(\epsilon,p)$ from the origin [36, 37, 38]. [^5] Other properties of DSR models, such as the laws of addition of energy and momentum, define torsion and nonmetricity [37, 41] Because of the relativistic properties of DSR models, these possibly curved momentum space metrics are maximally symmetric, that is they allow for a full set of relativistic generators. For instance, the momentum space metric of the DSR model we are considering here is known to be that of a de Sitter manifold [39, 40, 37, 38, 41, 42, 43]: the mass-shell relation (7) can be obtained from the metric $\tilde{g}(\epsilon)=\text{diag}(1,-e^{2\ell\epsilon})$, which is indeed a de Sitter metric on momentum space. [^6] An example of anti-de Sitter momentum spaces can be found in [44].
@@ VI Time Dilation
In special relativity, the proper time interval $\Delta\tau$ of a particle and the corresponding time interval $\Delta t$ measured by a boosted observer are related by $\Delta t=\gamma_{sr}\Delta\tau$, where the time dilation factor is $\gamma_{sr}$$\cosh{\eta}=\epsilon/M$, written in terms of rapidity $\eta$ or of energy $\epsilon$ of the particle, respectively. This special-relativistic result can be equivalently obtained either from the Lorentz transformation of the time interval or from the Minkowski metric. In this Section, we use the relations derived so far to compute the time dilation factor in the DSR model, showing that the two methods give equivalent (but deformed) results also in the DSR scenario.
@@
where we defined the coordinate velocity $V\equivdx$/dtdx/dt as in Section IV and we used the fact that $ds^{\prime}$$d\tau$ in the particle’s rest frame. Using (15) and (31) with $\eta^{\prime}=-\eta$ to write $\epsilon$ and $V$ in terms of the rest-frame quantities $(\epsilon^{\prime}$$M,p^{\prime}$==$0)$, we can integrate Eq. (38) to find the same relation as in Eq. (37). Therefore, in our DSR model one has $\Delta t=\gamma_{dsr}\Delta\tau$, with
@@
The time dilation factor $\gamma_{dsr}$ reduces to the standard factor $\gamma_{sr}$$\cosh{\eta}=\epsilon/M$ in the undeformed limit $\ell\to 0$ and is fully analytical in the deformation parameter $\ell$, in contrast to previous proposals described in Section I. Moreover, $\gamma_{dsr}$ is well-defined, positive and monotonous for all possible values of $M$ and $\epsilon$ (or $\eta$), that is, it has no pathologies and always preserves causality. Looking at the infinite-boost limit, when $\ell>0$ the time dilation factor $\gamma_{dsr}$ saturates to the finite value $\coth{(\ell M)}$ for $\eta\to\pm\infty$, while for negative deformation parameter $\ell<0$ the infinite boost limit corresponds to $\eta\to\eta_{\pm}$ and the factor $\gamma_{dsr}$ is divergent. [^8] The infinite-boost limit can be equivalently taken in terms of energy: for both signs of $\ell$ it reads $\epsilon\to\infty$ and one finds the same saturation property of $\gamma_{dsr}$, as can be directly checked from (41). This can be compared to special relativity, where the time dilation factor $\gamma_{sr}$ diverges for $\eta\to\pm\infty$.
@@ VII Summary and Outlook
In this paper, we investigated the time dilation effect within theories characterized by a Planck-scale modified dispersion relation of the form (1). Having observed that previous proposals, described by Eq. (2) and derived independently in [13] and [15], are consistent neither with standard Lorentz-invariance violation scenarios nor with DSR scenarios, we derived the time dilation effect within a fully DSR-relativistic model compatible with the dispersion relation (1), and adopting standard (i.e. trivial) Poisson brackets among spacetime coordinates.
We derived the finite boost transformations on all phase-space variables and to all orders in the deformation parameter. We showed that free particle worldlines are fully covariant under these boost transformations, generalizing previous results obtained at first order in the deformation parameter, and we identified the covariant spacetime metric. These results provide the ingredients needed to compute the time dilation effect, encoded in Eq.(40). Having obtained a formula valid to all orders in the deformation parameter, we could verify that the associated dilation factor $\gamma_{dsr}$ is well-defined and monotonous for all boosts. This is in contrast with other descriptions of time dilation beyond special relativity, encoded in Eq. (2), as discussed in Section I. This qualitative difference translates into different phenomenological implications.
@@
Not surprisingly, the correction with respect to the standard factor is governed by $\ell\epsilon$, a quantity that is small as long as the energy is subplanckian. In other words, our fully relativistic DSR derivation provides an analytical deformation of the standard time dilation formula. Because of this, phenomenological effects that were estimated to be significant for the scenarios leading to Eq. (2) are much smaller in this scenario (see Fig. 1 and Fig. 2).
Figure 1: Time dilation factors as a function of energy for high-energy muons ($\epsilon\gg M\simeq 100\,\text{MeV}$). The deformation scale is taken to be Planckian ($\ell^{-1}\simeq 10^{19}\,\text{GeV}$). The heuristic LIV prediction, Eq. (2), and the Finsler prediction, Eq. (2) with opposite sign, are compared to our DSR prediction in Eq. (41) and to the standard special relativity (SR) result, Eq. (42). Both the LIV and Finsler curves deviate significantly from the special-relativistic curve already at relatively low energies around $10^{6}\,\text{GeV}$. In contrast, the DSR curve essentially overlaps with the special-relativistic one throughout this low-energy range.
Figure 2: Time dilation factors as a function of energy for high-energy muons ($\epsilon\gg M\simeq 100\,\text{MeV}$). As in Fig. 1, we take $\ell^{-1}\simeq 10^{19}\,\text{GeV}$. The DSR prediction in Eq. (41) starts to deviate significantly from the special-relativistic (SR) one only at high energies around $10^{18}\,\text{GeV}$.
@@
Evidently, a natural next step for this research direction will be to verify whether the features we found in the specific DSR scenario here adopted are also present in other DSR scenarios. We conjecture that most of our results will be of general applicability, since in deriving them we developed the intuition that we were dealing with features connected primarily with just the presence of relativistic invariance, rather than with specific details of the DSR scenario.